OpenStudy (anonymous):

Solve 3 square root of 45 x cubed end root plus 2 square root of 12 end root plus square root of 27 end root minus 3 square root of 20 x cubed Okay, I understand what you do when there are like terms but there isn't any in this one

OpenStudy (amistre64):

you might have heard of this fancy thing called notation ....

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

you might want to post this in the writing section for a suitable translation :)

OpenStudy (anonymous):

But this is a math problem

OpenStudy (amistre64):

realy? i couldnt tell let me see if i can rewrite it

OpenStudy (amistre64):

3 sqrt(45x^3) + 2 sqrt(12) +sqrt(27) -3 sqrt(20x^3)

OpenStudy (anonymous):

Yes that is how it is written but idk how to solve when there are not any like terms

OpenStudy (amistre64):

you reduce those sqrts and see if they got any likes about them

OpenStudy (anonymous):

How do I reduce?

OpenStudy (amistre64):

sqrt(45) for instance; does 45 have a perfect square factor?

OpenStudy (anonymous):

Uh I don't think so

OpenStudy (amistre64):

how would we check to see if it does instead of just guessing at it?

OpenStudy (anonymous):

factor it out?

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

what are the factors of 45

OpenStudy (anonymous):

9 and 5

OpenStudy (amistre64):

those lok nice; soo, sqrt(9*5) = sqrt(9) * sqrt(5) do either of these reduce?

OpenStudy (anonymous):

yes the 9

OpenStudy (anonymous):

into 3

OpenStudy (amistre64):

good :) and does sqrt(x^3) reduce?

OpenStudy (anonymous):

nope

OpenStudy (amistre64):

lol, you sure? sqrt(x^2*x) = sqrt(x^2) * sqrt(x) might have a perfect square about it

OpenStudy (anonymous):

Okay so it does..lol

OpenStudy (amistre64):

this might take awhile ...

OpenStudy (amistre64):

sooo 3 sqrt(45x^3) = 9sqrt(5x) we agree?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

3 sqrt(45x^3) + 2 sqrt(12) +sqrt(27) -3 sqrt(20x^3) 9 sqrt(5x) + 2 sqrt(12) +sqrt(27) -3 sqrt(20x^3) that 20x^3 looks like it might be a "like term" for our 45x^3 fiasco

OpenStudy (amistre64):

does 20x^3 have any perfect square factors that we can reduce/simplify ?

OpenStudy (anonymous):

4

OpenStudy (amistre64):

4 and x^2 to make life a little quicker :)

OpenStudy (amistre64):

9x sqrt(5x) + 2 sqrt(12) +sqrt(27) -6x sqrt(5x) now we have "like terms" the middle 2 wil work out in the same fashion

OpenStudy (anonymous):

so the first half is 3x sqrt(5x) right?

OpenStudy (amistre64):

yes and a bit of trivia: "like terms" is a bad name for it, since the math of it calls them a greatest common factor

OpenStudy (anonymous):

Okay and for the second half we have to factor out 12 and 27?

OpenStudy (amistre64):

yes; into 4*3 and 9*3

OpenStudy (anonymous):

so what do you do after that?

OpenStudy (amistre64):

nothing different from what we did with the first 2 really

OpenStudy (amistre64):

factor out the perfect squares and simplify and comdine

OpenStudy (anonymous):

so its 7 sqrt 3?

OpenStudy (amistre64):

+ 4 sqrt(3) +3sqrt(3) yep, good job ;)

OpenStudy (anonymous):

Thanks you so much, I'm sorry I'm just really bad at math

OpenStudy (amistre64):

it takes practice, dedication, and six fingers on one hand ...